Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day11


Crew 99 Log Book for February 8, 2011

Photos of the day





Commander's Check-In Report

 Donna Viola Reporting

 Crew Physical Status: Good


 Time Departed/Returned from EVA: EVA #17, 11:45-12:30; EVA #18,

 12:45-13:15; EVA #19, 15:00-17:15


 Brief Narrative of Field Mission Results: EVA planner, scoping out

 sites for geologic analysis tomorrow/Thursday


 EVA Data/Interpretations: as per EVA reports


 Engineering/Hab Maintenance: Nothing major to report.


 Report Transmission Schedule: Engineering, EVA#17, EVA#18, EVA#19,

 Commander's, Chef's, Science. By 9:00pm


 Plans for Tomorrow: Additional mission planner EVAs; Endolith sample 

collection; Collecting geological data


 Inventory: N/a


 Support Requested: None


 Miscellaneous: Happy birthday to crew engineer/astronomer, Max!


Commander's Report

Donna Viola Reporting

Today, we realized just how little time we have remaining here at MDRS. This morning, we held a team meeting to figure out all the objectives we'd like to complete over our last few days. Andie decided to run two mission planner EVAs, working on planning contingencies both in the field and remotely, in communication with someone back at the Hab.


We made sandwiches for lunch, using bread prepared last night that didn't look especially attractive, but still tasted good! After that, Heidi, Kevin, and I took the ATVs to scope out some sites for geological study and for endolith and soil sample collection. Upon our return, Max and Sukrit were busy preparing dinner - pasta with homemade sauce!


Today was also Max's birthday, so Andie baked a chocolate cake based on a recipe we found in the Hab. We interrupted a conversation about observing the different layers in an exoplanet's atmosphere to sing Happy Birthday - then Max and Sukrit jumped right back into their discussion without missing a beat. The astronomers have since ventured out to the observatory, where they hope to polar align the telescope tonight.


Engineering Report

Kevin Newman Reporting


Generator/Electricity:


· Kitty (Diesel Generator) (Used? Yes/No): Yes


· Honey (Gas Generator) (Used? Yes/No): No


· Battery Charge Level: Checked by DG


o 10.59 V everstart


o 12.37 V armor plate


o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run


Fuel Status:


· Diesel (Full/0.75/0.5/0.25/Empty): Full


· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.72


· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 16 Gallons


Notes/Comments: Based on the random spawning of molecules in the diesel tank, we have surely discovered spontaneous generation and will soon be writing a paper in the name of MDRS. Intelligent design is also making a comeback.


Oil:


· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG


· Honey Oil Quantity (# of Quarts in storage): Leave it to DG


· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench


· Car Oil Quantity (# of Quarts in storage):


Water Status:


· Outside Potable Water Tank Level (inches from bottom): 17


· Trailer Potable Water Tank Level (inches from bottom): 5/6 Full


· Hab Potable Water Tank Level (inches from bottom): 6


· Potable Water Meter Reading: 418223



Notes/Comments:


GreenHab:


· Condition of plants in Tank 1: Nothing


· Condition of plants in Tank 2: Ok


· Condition of plants in Tank 3: Good


· Greenhab notes/ comments:


Observatory:


· Telescope Used? (Yes/no) Yes


Observatory Notes: Hopefully will be polar-aligned by the end of tonight.


Transportation:


Hab Car (NOT RENTAL)


· Used (yes/no): No


· Oil Checked (yes/no): No


· Tire Status: Not checked


· Odometer at end of day: 123611.6


· Notes/Comments on Hab Car:


ATV 1 (Opportunity):


· Used: Yes


· Oil Checked: Yes


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


ATV 2 (Spirit):


· Used: Yes


· Oil Checked: Yes


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


ATV 3 (Viking I):


· Used: No


· Oil Checked: No


· Fuel Consumed: None


· Tire Status: Not Checked


· Comments:


ATV 4 (Viking II):


· Used: Yes


· Oil Checked: Yes


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


Heating and Ventilation:


· Thermostat setting upstairs (Farenheit): 66


· Actual temperature upstairs (Farenheit): 71


· Thermostat setting downstairs (Farenheit): 65


· Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:


Notes/comments on internet/computers?: Behaving well


General Engineering

1. Notes/Comments: Turning down the propane heater downstairs seems to have dropped the Hab temperature dramatically. The heater has been temporarily turned up again because we were a little cold 


Clear Skies!


Science Report LAMBDA

Donna Viola Reporting

A few days ago, the team began a trial of our LAMBDA experiment, where we mixed a soil sample with water and placed it in a microbial fuel cell hoping that we will be able to detect microbial life by measuring the current generated by the soil sample. For the last few days, we have been collecting voltage data.


Tomorrow, we plan on beginning our control run, where we will sterilize our sample and collect more voltage data. If the voltage drops after sterilization, it will imply that the higher voltage in the first run was due to metabolizing microbes. We are also planning on confirming any results in the lab post-rotation using fluorescent microscopy.



Science Report (Geology)

Heidi Beemer Reporting

EVA#: 3, 7, 8

Crew Members Involved: Heidi Beemer

Site 1 Location: N 38˚ 24.141 W 110˚ 47.556

Site 2 Location: To be Determined

Site 3 Location: To be Determined

Transit Mode: ATV/ Walking / ATV

Objectives:  Due to radiation being present on Mars, settlers will most likely seek refuge under the earth. At first caves will be used, however, as civilizations grow and caves become filled, tunneling will be the next logical step. An important factor in any tunneling project is the roof structure holding up the tunnel. It is imperative that the rock structure is analyzed before a tunneling project begins. The easiest way to classify the geomechanics of a site is to use a rock mass rating or RMR (Bieniawski, 1972). This rating uses six parameters to classify rock formations. These parameters are: uniaxial compressive strength of rock material, Rock Quality Designation (RQD), spacing of discontinuities, condition of discontinuities, groundwater conditions, and orientation of discontinuities. These six parameters are given number ratings based on the conditions for each. The final RMR number is a number 1-100 and is the total of all the ratings combined. These parameters are based on field reports and the rock mass rating will tell the engineer the behavior of the rock, providing quantitative data for engineering design.

Due to the unique geology of “Mars” in the San Rafael Swell, there are many locations in which tunnel type shelters can be built for expansion during colonization. The purpose of this investigation is to calculate a RMR rating of good rock or better (61-100) for three possible shelter locations.

Results: A RMR rating of 63 was determined from the first site. This was determined on the following parameters:

Uniaxial compressive strength of rock material: Medium High 50-100 MPa......13

Rock Quality Designation (RQD): Fair Quality 70%......................................7                        

Spacing of discontinuities: Moderate 200-600mm.........................................10

Condition of discontinuities: Moderately weathered........................................20

Groundwater conditions: Completely dry......................................................15

Orientation of discontinuities: Drive w/ dip Favorable......................................-2

                                                                                                         Total: 63

This rating is a class II with a description of Good Rock. Once tunneling begins an adjusted RMR scale is used to assess conditions once blasting has occurred and stress is induced. RMR can also be used to calculate stand up time for the tunnel and deciding the support needed.

This is based on case histories of previous occurrences of roof falls.

Two more locations on " Mars" will be assessed for geomechanic classification by the end of the week.


Astronomy Report

Max Fagin Rerporting

Participants: Sukrit Ranjan and Max Fagin.


Time: 22:00-23:30


Objectives: Identify the source of error in all previous polar alignments, and align the telescope mount to the CNP if needed.


Results: The telescope was powered up successfully, but as soon as we began to slew, the black UPS under the table began emitting a shrill tone, and the webcam computer lost power.  Power to the rest of the observatory, including the telescope, was not interrupted.  We reset the UPS, and restarted the computer.  We then communicated with the Hab to confirm that the webcams were back online.  While waiting for the computer to restart, we experimented with controlling the CCD via the Equinox Pro Observatory program.  We were not able to establish a connection with the CCD via this program.


We then attempted to align the telescope multiple times, using various combinations of guide stars.  Three separate two star alignments produced equally poor results, with accumulated errors of several degrees.  We concluded that the error most likely lay in the mount itself being off level. Progress during these aligns was impeded by the hand controller's tendency to revert to greek characters if it got to cold.  Warming the controller in our ungloved hands fixed this problem each time, but it did leave us with very cold hands.


The bubble level on the mount was woefully inadequate for our purposes. As such, we used the theodolite function on the iPhone to level the mount. The mount was found to be tipped to the North by ~.5°.  We corrected this tilt by adjusting the bolts at the base of the mount.


We originally planned to align the mount by using the telescope's built in "Polar Align" feature.  However, as we had been warned, the menu tree shown in the Celestron Manual contained many entries which were not available on the hand mount itself.  The Polar Align feature was one of them.


We thus attempted to perform a manual polar alignment.  We ordered the telescope to DEC +90°, and while peering through the eye piece, began slewing rapidly in RA.  If the telescope were truly at DEC +90°, the star field visible through the eyepiece would rotate about a point in the center of the FOV as we slewed through RA.  If the alt-az setting on the mount were correct, this point of rotation would be located on the celestial north pole. By making fine adjustments in DEC while slewing rapidly through RA, we were able to point the telescope along it's true axis of rotation.  The axis was found to deviate by ~3° from the CNP.


By incremental adjustments of the altitude and azimuth of the mount (via the hex wrench and thumb screws) and by keeping the telescope at DEC +90°, we were able to center the telescope's axis of rotation ~.5° to the east of Polaris, where the true CNP was located at the time.  After making these adjustments, slewing the telescope in RA caused the star field visible in the eye piece to rotate about the CNP, which was itself located in the center of the FOV.  At this point however, the cold became unbearable, and we turned in for the night.  We will check tomorrow to see if the telescope is now truly polarly aligned.


During shutdown, the UPS box again began emitting a shrill tone, and the observatory computer again lost power.  We restarted both the UPS and computer, and closed out the observatory.


Requests for Assistance:


1) Explanation why the power supply box failed, and steps we can take to avoid it's failure in the future.


2) Explanation as to why the menu tree on the hand controller deviates so drastically from the menu tree in the Celestron manual.


EVA 17 and 18 Report

Andie Gilkey Reporting

Time: 12-12:30; 13-13:30

Crew Members:

EVA 17: Andie Gilkey, Heidi Beemer
EVA 18: Andie Gilkey, Max Fagin

Site Location: Thielges Flats, UTM 518500 4251100


Transit Mode: Foot

Objectives: Use EVA mission planner to plan mission and contingency while in the field

Results: Mission planner was successfully used to find the site and re-plan in the field

Lessons Learned: It is difficult to see the iPad screen in the bright sunlight


EVA 19 Report

Kevin Newman Reporting

Time: 15:15-17:15

Crew Members Heidi Beemer, Kevin Newman, Donna Viola

Site Location: Unnamed ATV trail, ending at N 38 24.344' W 110 45.161' 

Transit Mode: ATV

Objectives: Identify geological sites for analysis tomorrow

Results: Several sites found and photographed along unnamed ATV trail (will be added to database of trails and waypoints). We also think we found an ATV trail to Hanksville which was mentioned by a local.

Lessons Learned: Keep your ATV key on you at all times. If you go down a hill, first make sure that you can get back up! The most beautifully scenic locations are often the most difficult to get to, and the terrain in certain areas can consists of deep loose sand that makes steep ascension difficult.


Journalist Report

Sukrit Ranjan Reporting

Today the team took stock of their progress and started to really push in order to achieve all science objectives by rotation end. We deployed back-to-back EVAs all day long to gather data, and spent 1.5 hours in the freezing observatory realigning the telescope. At dinner, we celebrated Max's birthday with home-made pasta sauce and cake. It was delicious. Onward!


Cook's Report

Sukrit Ranjan Reporting

1a) Today was

-A cooking day

1b) Was today a special day celebrated at the main meal with special

food or activity?

Yes; today we celebrated Max's birthday.

2)  List the foods served at today's main meal (usually the evening

meal), giving full names of each.

Pasta with tomato sauce with vegetables

Chocolate cake

3) List any main meal foods not finished by the crew (leftovers stored

for later use or discarded)  (Example:  2 servings potato soup stored;

2 servings rice burned/inedible, discarded)

-6 servings of cake left

-2 servings of pasta sauce left

4)  Recipes for anything more complicated than following package

directions exactly, or rehydrating.

- Followed recipe in hab recipe collection for chocolate cake.

- Followed recipe in hab recipe collection for pasta sauce

  (roughly: 3 handfuls mushroom, onion, 1 handful tomato bits, 1 handful spinach, tomato powder, herbs to taste; in our case, a bit of ginger, lots of oregano, a moderate amount of basil, a bit of ground red pepper, a bit of salt, lots of garlic salt)

-people added tofu or rehydrated beef bits to taste

5)  Comments/questions for Kim and Jean, the food study investigators

- Cake baked unevenly, reflected temperature gradient in oven.